### v2

@@ C Description of the Identified Sparse Factors
20090112 -- By concerns about Citigroup’s potential deal with Morgan Stanley - and the start of the fourth-quarter earnings reporting period Company-specific factors
@@
20090114 -- As a bleak retail sales report and more dour news from the banking sector amplified fears of a prolonged recession Economic indicators
@@
20090120 -- As investors looked beyond President Barack Obama’s historic inauguration to the battered economy he inherits Market sentiment
@@
20090121 $+$ As investors welcomed IBM’s earnings and scooped up bank shares hit hard in the recent retreat Company-specific factors
@@
20090122 -- As a management shakeup at Bank of America and Microsoft’s earnings disappointment weighed on investor sentiment Company-specific factors
@@
20090128 $+$ As investors took comfort in reports that the Obama administration and the Federal Reserve are taking steps to get credit flowing again and help staunch the economic slowdown Government policies
@@
20090129 -- Following more dire news on earnings, housing and employment Economic indicators
@@
20090130 -- As investors eyed abysmal reports on economic growth and quarterly earnings Economic indicators
@@
20090206 $+$ As optimism about the government’s economic stimulus bill and the new version of the bank bailout plan countered unease following the brutal January jobs report Government policies, Economic indicators
@@
20090210 -- As the government’s bank rescue plan failed to reassure investors burned by the 14-month old recession Government policies
@@
20090217 -- On fears that the government’s efforts to slow the recession won’t be sufficient Market sentiment
@@
20090219 -- As fears of a prolonged recession sent stock investors heading for the exits Market sentiment
@@
20090223 -- As investors continue to worry that the government’s efforts to slow the recession won’t be sufficient Market sentiment
@@
20090224 $+$ After comments from Fed Chairman Ben Bernanke that downplayed bank takeover fears helped to spark a big rally Government policies, Market sentiment
@@
20090227 -- On worries about the government taking a bigger chunk of Citigroup and a bleak reading on the economy, again touching 12-year lows Government policies, Economic indicators
@@
20090302 -- After insurance company American International Group’s huge quarterly loss added to worries about the financial sector and the economy Company-specific factors
@@
20090304 $+$ Following reports that China’s economy may be improving and as government officials unveiled details of the \$$75$ billion foreclosure fix China
@@
20090305 -- As investors waded through more grim news: GM said its survival is in doubt, bank shares took a beating, and Citigroup fell below a buck Company-specific factors
@@
20090310 $+$ After Citigroup cooled some worries about its future and regulators said they may reinstate a key trading rule Company-specific factors, Government policies
@@
20090312 $+$ As investors scooped up banks and other shares hit in a selloff that left the Dow at 12-year lows Market sentiment
20090317 $+$ As investors continued to dig out from 12-year lows Market sentiment
@@
20090318 $+$ After the Federal Reserve said it would buy up to \$300 billion in long-term government bonds Government policies
20090320 -- As investors pulled back after the recent run Market sentiment
@@
20090323 $+$ After Treasury’s plan to buy up billions in bad bank assets and a better-than-expected existing home sales report raised hopes that the economy is stabilizing Government policies, Economic indicators
20090326 $+$ As the March market run shows its legs Market sentiment
@@
20090330 -- As auto and bank woes spark a selloff after the rally Economic indicators
@@
20090402 $+$ After key accounting rule that has impact on banks is changed G-20 also in focus Government policies
@@
20090407 -- After a four-week advance, on worries about banks and autos and the start of the quarterly reporting period Market sentiment
@@
20090409 $+$ After Wells Fargo forecast a nearly \$$3$ billion quarterly profit, adding to hopes that the banking sector is stabilizing Company-specific factors
@@
20090414 -- After a weaker-than-expected retail sales report gave investors a reason to retreat following a five-week run Company-specific factors, Market sentiment
@@
20090416 $+$ The major stock gauges touch the best levels in months on JPMorgan Chase earnings Company-specific factors
@@
20090420 -- On worries about financial sector earnings, despite Bank of America’s better-than-expected quarterly results Market sentiment
@@
20090421 $+$ As worries about corporate results were countered by renewed hopes that the financial sector is closer to stabilizing Market sentiment
@@
20090424 $+$ After Ford, Microsoft and American Express reported results that met or topped analysts’ expectations Company-specific factors
@@
20090429 $+$ After the Federal Reserve held interest rates steady, as expected, but issued a slightly more upbeat economic outlook Government policies
@@
20090504 $+$ As a better-than-expected housing market report intensified hopes that the economy is closer to stabilizing Economic indicators
@@
20090508 $+$ After a government report showed employers cut fewer jobs than expected last month Economic indicators
@@
20090511 -- As investors took a step back after propelling the major stock gauges by more than $30\%$ each in just two months Market sentiment
@@
20090513 -- After a weaker-than-expected retail sales report gave investors a reason to retreat Economic indicators
@@
20090518 $+$ After positive news about the US housing market, including an upbeat profit forecast from Lowes, as well as an upgrade of Bank of America Economic indicators, Company-specific factors
@@
20090526 $+$ After a report showing consumer confidence hit an eight-month high offset dismal housing news Market sentiment
@@
20090601 $+$ As better-than-expected readings on manufacturing activity raised hopes that a global economic recovery is brewing Economic indicators
@@
20090615 -- As weaker oil prices and more geopolitical unrest raised worries that the recession may not be waning as soon as some had hoped Oil price, Global events (Geopolitical events)
@@
20090622 -- As the World Bank’s weak outlook on global growth and a selloff in commodity prices sent investors heading for the exits Market sentiment
@@
20090702 -- After a worse-than-expected jobs report hammered hopes that the economy is close to stabilizing Economic indicators
@@
20090713 $+$ As investors welcomed an analysts’ improved outlook on Goldman Sachs one day ahead of its quarterly report Market sentiment, Company-specific factors
@@
20090715 $+$ After Intel’s forecast for a second-half pickup and the Federal Reserve’s improved outlook reassured wary investors Company-specific factors
@@
20090723 $+$ As investors welcomed better-than-expected quarterly results and home sales Economic indicators
@@
20090817 -- As worries that nervous consumers will pressure a fragile recovery dragged stocks lower after a five-month advance Market sentiment
@@
20090901 -- On worries that the market gains have raced ahead of any economic recovery Market sentiment
@@
20091001 -- After a bigger-than-expected rise in weekly jobless claims and a weaker-than-expected reading on manufacturing sparked worries about the pace of the economic recovery Economic indicators
@@
20091028 -- As a weaker-than-expected new home sales report added to questions about the strength of the economic recovery Economic indicators
@@
20091029 $+$ As a strong report on economic growth in the third quarter reassured investors that the recovery is on track Economic indicators
@@
20091030 -- As investors dumped a variety of shares at the end of a rough week and choppy month on Wall Street Market sentiment
20100204 -- A growing debt crisis in Europe Europe (Debt crisis)
@@
20100427 -- After Standard & Poors cut Greece’s debt rating to junk and lowered Portugal’s debt rating, Europe (Debt crisis)
@@
20100504 -- On worries that the global recovery could suffer if Europe’s efforts to contain Greece’s debt problems don’t succeed, and if China’s efforts to slow its booming economy go too far Europe (Debt crisis), China
20100506 -- Fears about the spread of the European debt crisis Europe (Debt crisis)
@@
20100510 $+$ After European officials approved a nearly $1$ trillion rescue plan to contain the debt crisis in troubled nations and stabilize the euro Government policies
@@
20100520 -- Worries about how the European debt crisis and slump in the euro will impact the global recovery fueled the selling Europe (Debt crisis)
@@
20100527 $+$ After Chinese officials dismissed reports that they’re reviewing their nation’s investment in European bonds amid concerns about the continent’s debt problems China
@@
20100601 -- As worries about the global economic outlook overshadowed better-than-expected readings on the US economy Market sentiment
@@
20100604 -- After a government report showed employers added fewer jobs than expected last month and the euro plunged to a new 4-year low, reviving worries about the health of the European economyEconomic Indicators Economic indicators, Europe (Debt crisis)
@@
20100610 $+$ As concerns over Europe’s debt crisis and its impact on the global recovery were calmed by a sharp boost in Chinese exports and a strengthening euro Europe (Debt crisis), China
20100629 -- After a big drop in consumer confidence and signs of a bigger slowdown in the global economy Market sentiment
20100707 $+$ As investors came back after the recent bloodletting, spurred on by State Street’s improved earnings forecast Economic indicators
20100716 -- After financial firms Bank of America and Citigroup reported weaker quarterly revenue and a plunge in consumer sentiment revived concerns about the economic outlook Company-specific factors
20100722 $+$ After better-than-expected earnings and forecasts from 3M, Caterpillar, AT&T and UPS helped reassure investors about the pace of the economic recovery Economic indicators, Company-specific factors
20100811 -- After a report showed the US trade gap widened, and foreign data cast doubt on overseas demand for American goods Economic indicators
20100901 $+$ As investors cheered signs of strength in the manufacturing sector Economic indicators
$2$- Libya’s escalating political crisis sparked a sharp sell-off in US stocks, as oil prices continued to skyrocket Global events (geopolitical events)
20110601 -- Weak economic data has started to snowball Economic indicators
20110802 -- As fears about a weak US economy were enflamed after investors got another disappointing economic report - this time on consumer spending Economic indicators
20110804 -- As fear about the global economy spooked investors Market sentiment
20110808 -- As the debt crisis in Europe, lackluster economic news and a downgrade to the US credit rating Europe (Debt crisis), Credit risk
20110809 $+$ After the Federal Reserve said it will keep interest rates exceptionally low until 2013 Government policies
20110810 -- As investors were confronted with mounting fears about Europe’s ongoing debt crisis, this time in France Europe (Debt crisis)
20110811 $+$ On positive earnings and labor market news Economic indicators
20110818 -- As renewed concerns about the US and global economies Market sentiment
20110823 $+$ Following a report from the FDIC that showed the number of failing banks shrank for the first time in nearly five years Economic indicators
20110829 $+$ A Greek bank deal, a solid US consumer spending report and relief that Hurricane Irene caused less damage than expected Government policies
20110902 -- After a government report showing no job growth in August stoked fears that the US may be headed into another recession Economic indicators
20110907 $+$ The gains came as concerns over Europe’s debt crisis eased and investors geared up for President Obama’s highly anticipated jobs speech Thursday evening Europe (Debt crisis)
20110909 -- As bad news out of Europe kept piling up Europe (Debt crisis)
20110921 -- Fed disappoints Government policies
20110922 -- Fear factor Market sentiment
20110928 -- Worries over stalled global growth prospects and concerns that European leaders may not be moving fast enough to solve the region’s debt problems Market sentiment
20110930 -- As investors remain worried about the debt crisis in Europe and the outlook for global economic growth Market sentiment
20111003 -- With worries about Greece’s solvency still in the spotlight Market sentiment
20111004 $+$ Bear market bounce Market sentiment
20111010 $+$ As investors cheered a pledge from European leaders to unveil a plan for solving the eurozone’s debt crisis by the end of the month Government policies
@@
20111017 -- As worries about Europe’s debt crisis dominated Market sentiment, Europe (Debt crisis)
@@
20111018 $+$ Following a report suggesting that Europe’s bailout fund may get a big boost Market sentiment, Government policies
@@
20111027 $+$ After European Union leaders agreed to expand Europe’s bailout fund and take major losses on Greek bonds Government policies
@@
20111031 -- As investors continued to scrutinize the eurozone debt deal Europe (Debt crisis)
@@
20111101 -- New fears about the fate of the European rescue plan Market sentiment
20111109 -- Italy fears Europe (Debt crisis)
@@
20111123 -- As eurozone fears rumbled on and a preliminary report showed that Chinese manufacturing slowed sharply Europe (Debt crisis), China
20111128 $+$ On robust Black Friday sales Economic indicators
@@
20111130 $+$ After the Federal Reserve said it will work with other central banks to support the global economy Government policies
@@
20111220 $+$ As concerns about the European debt crisis eased and investors welcomed signs of strength in the US housing market Market sentiment, Economic indicators
@@
20120601 -- As Europe’s debt crisis remains unresolved and the US economy is showing new signs of distress Europe (Debt crisis), Market sentiment
@@
20120629 $+$ A deal among European leaders to help struggling eurozone banks buoyed global markets Government policies
@@
20130415 -- Following the news of explosions at the Boston Marathon Global events (terrorist attack)
20150824 -- Deep fears about China’s economic slowdown China
@@
20150826 $+$ The huge rally represents a rebound following six days of dramatic selling that was driven by serious concerns about how China’s slowing economy will impact the rest of the world China
20150901 -- Following more fears about a slowdown in China’s economy China
20160113 -- Low oil prices, as concerns about global economic slowdown Oil price
@@
20160129 $+$ A realization that if the US avoids a recession – as most economists think it will – beaten-down stocks could be a good buy Market sentiment
@@
20160624 -- British voters chose to leave the European Union Europe (Brexit)
@@
20160627 -- British voters chose to leave the European Union Europe (Brexit)

### v3

@@ D Description of the Identified Sparse Factors
20090112 - By concerns about Citigroup’s potential deal with Morgan Stanley - and the start of the fourth-quarter earnings reporting period Company-specific factors
@@
20090114 - As a bleak retail sales report and more dour news from the banking sector amplified fears of a prolonged recession Economic indicators
@@
20090120 - As investors looked beyond President Barack Obama’s historic inauguration to the battered economy he inherits Market sentiment
@@
20090121 ++ As investors welcomed IBM’s earnings and scooped up bank shares hit hard in the recent retreat Company-specific factors
@@
20090122 - As a management shakeup at Bank of America and Microsoft’s earnings disappointment weighed on investor sentiment Company-specific factors
@@
20090128 ++ As investors took comfort in reports that the Obama administration and the Federal Reserve are taking steps to get credit flowing again and help staunch the economic slowdown Government policies
@@
20090129 - Following more dire news on earnings, housing and employment Economic indicators
@@
20090130 - As investors eyed abysmal reports on economic growth and quarterly earnings Economic indicators
@@
20090206 ++ As optimism about the government’s economic stimulus bill and the new version of the bank bailout plan countered unease following the brutal January jobs report Government policies, Economic indicators
@@
20090210 - As the government’s bank rescue plan failed to reassure investors burned by the 14-month old recession Government policies
@@
20090217 - On fears that the government’s efforts to slow the recession won’t be sufficient Market sentiment
@@
20090219 - As fears of a prolonged recession sent stock investors heading for the exits Market sentiment
@@
20090223 - As investors continue to worry that the government’s efforts to slow the recession won’t be sufficient Market sentiment
@@
20090224 ++ After comments from Fed Chairman Ben Bernanke that downplayed bank takeover fears helped to spark a big rally Government policies, Market sentiment
@@
20090227 - On worries about the government taking a bigger chunk of Citigroup and a bleak reading on the economy, again touching 12-year lows Government policies, Economic indicators
@@
20090302 - After insurance company American International Group’s huge quarterly loss added to worries about the financial sector and the economy Company-specific factors
@@
20090304 ++ Following reports that China’s economy may be improving and as government officials unveiled details of the \$$75$ billion foreclosure fix China
@@
20090305 - As investors waded through more grim news: GM said its survival is in doubt, bank shares took a beating, and Citigroup fell below a buck Company-specific factors
@@
20090310 ++ After Citigroup cooled some worries about its future and regulators said they may reinstate a key trading rule Company-specific factors, Government policies
@@
20090312 ++ As investors scooped up banks and other shares hit in a selloff that left the Dow at 12-year lows Market sentiment
20090317 ++ As investors continued to dig out from 12-year lows Market sentiment
@@
20090318 ++ After the Federal Reserve said it would buy up to \$300 billion in long-term government bonds Government policies
20090320 - As investors pulled back after the recent run Market sentiment
@@
20090323 ++ After Treasury’s plan to buy up billions in bad bank assets and a better-than-expected existing home sales report raised hopes that the economy is stabilizing Government policies, Economic indicators
20090326 ++ As the March market run shows its legs Market sentiment
@@
20090330 - As auto and bank woes spark a selloff after the rally Economic indicators
@@
20090402 ++ After key accounting rule that has impact on banks is changed G-20 also in focus Government policies
@@
20090407 - After a four-week advance, on worries about banks and autos and the start of the quarterly reporting period Market sentiment
@@
20090409 ++ After Wells Fargo forecast a nearly \$33 billion quarterly profit, adding to hopes that the banking sector is stabilizing Company-specific factors
@@
20090414 - After a weaker-than-expected retail sales report gave investors a reason to retreat following a five-week run Company-specific factors, Market sentiment
@@
20090416 ++ The major stock gauges touch the best levels in months on JPMorgan Chase earnings Company-specific factors
@@
20090420 - On worries about financial sector earnings, despite Bank of America’s better-than-expected quarterly results Market sentiment
@@
20090421 ++ As worries about corporate results were countered by renewed hopes that the financial sector is closer to stabilizing Market sentiment
@@
20090424 ++ After Ford, Microsoft and American Express reported results that met or topped analysts’ expectations Company-specific factors
@@
20090429 ++ After the Federal Reserve held interest rates steady, as expected, but issued a slightly more upbeat economic outlook Government policies
@@
20090504 ++ As a better-than-expected housing market report intensified hopes that the economy is closer to stabilizing Economic indicators
@@
20090508 ++ After a government report showed employers cut fewer jobs than expected last month Economic indicators
@@
20090511 - As investors took a step back after propelling the major stock gauges by more than $30\%$ each in just two months Market sentiment
@@
20090513 - After a weaker-than-expected retail sales report gave investors a reason to retreat Economic indicators
@@
20090518 ++ After positive news about the US housing market, including an upbeat profit forecast from Lowes, as well as an upgrade of Bank of America Economic indicators, Company-specific factors
@@
20090526 ++ After a report showing consumer confidence hit an eight-month high offset dismal housing news Market sentiment
@@
20090601 ++ As better-than-expected readings on manufacturing activity raised hopes that a global economic recovery is brewing Economic indicators
@@
20090615 - As weaker oil prices and more geopolitical unrest raised worries that the recession may not be waning as soon as some had hoped Oil price, Global events (Geopolitical events)
@@
20090622 - As the World Bank’s weak outlook on global growth and a selloff in commodity prices sent investors heading for the exits Market sentiment
@@
20090702 - After a worse-than-expected jobs report hammered hopes that the economy is close to stabilizing Economic indicators
@@
20090713 ++ As investors welcomed an analysts’ improved outlook on Goldman Sachs one day ahead of its quarterly report Market sentiment, Company-specific factors
@@
20090715 ++ After Intel’s forecast for a second-half pickup and the Federal Reserve’s improved outlook reassured wary investors Company-specific factors
@@
20090723 ++ As investors welcomed better-than-expected quarterly results and home sales Economic indicators
@@
20090817 - As worries that nervous consumers will pressure a fragile recovery dragged stocks lower after a five-month advance Market sentiment
@@
20090901 - On worries that the market gains have raced ahead of any economic recovery Market sentiment
@@
20091001 - After a bigger-than-expected rise in weekly jobless claims and a weaker-than-expected reading on manufacturing sparked worries about the pace of the economic recovery Economic indicators
@@
20091028 - As a weaker-than-expected new home sales report added to questions about the strength of the economic recovery Economic indicators
@@
20091029 ++ As a strong report on economic growth in the third quarter reassured investors that the recovery is on track Economic indicators
@@
20091030 - As investors dumped a variety of shares at the end of a rough week and choppy month on Wall Street Market sentiment
20091109 ++ As investor optimism gained momentum Market sentiment
20100204 - A growing debt crisis in Europe Europe (Debt crisis)
@@
20100427 - After Standard & Poors cut Greece’s debt rating to junk and lowered Portugal’s debt rating, Europe (Debt crisis)
@@
20100504 - On worries that the global recovery could suffer if Europe’s efforts to contain Greece’s debt problems don’t succeed, and if China’s efforts to slow its booming economy go too far Europe (Debt crisis), China
20100506 - Fears about the spread of the European debt crisis Europe (Debt crisis)
@@
20100510 ++ After European officials approved a nearly $1$ trillion rescue plan to contain the debt crisis in troubled nations and stabilize the euro Government policies
@@
20100520 - Worries about how the European debt crisis and slump in the euro will impact the global recovery fueled the selling Europe (Debt crisis)
@@
20100527 ++ After Chinese officials dismissed reports that they’re reviewing their nation’s investment in European bonds amid concerns about the continent’s debt problems China
@@
20100601 - As worries about the global economic outlook overshadowed better-than-expected readings on the US economy Market sentiment
20100602 ++ Fueled by a rebounding energy sector Company-specific factors
@@
20100604 - After a government report showed employers added fewer jobs than expected last month and the euro plunged to a new 4-year low, reviving worries about the health of the European economyEconomic Indicators Economic indicators, Europe (Debt crisis)
@@
20100610 ++ As concerns over Europe’s debt crisis and its impact on the global recovery were calmed by a sharp boost in Chinese exports and a strengthening euro Europe (Debt crisis), China
@@
20100615 ++ As worries about Europe’s debt woes hurting U.S. growth eased Europe (Debt crisis)
20100629 - After a big drop in consumer confidence and signs of a bigger slowdown in the global economy Market sentiment
20100707 ++ As investors came back after the recent bloodletting, spurred on by State Street’s improved earnings forecast Economic indicators
20100716 - After financial firms Bank of America and Citigroup reported weaker quarterly revenue and a plunge in consumer sentiment revived concerns about the economic outlook Company-specific factors
20100722 ++ After better-than-expected earnings and forecasts from 3M, Caterpillar, AT&T and UPS helped reassure investors about the pace of the economic recovery Economic indicators, Company-specific factors
20100811 - After a report showed the US trade gap widened, and foreign data cast doubt on overseas demand for American goods Economic indicators
20100901 ++ As investors cheered signs of strength in the manufacturing sector Economic indicators
20100924 ++ As a report on durable goods edged slightly ahead of forecasts Economic indicators
20110222 - Libya’s escalating political crisis sparked a sharp sell-off in US stocks, as oil prices continued to skyrocket Global events (geopolitical events)
20110601 - Weak economic data has started to snowball Economic indicators
20110727 - As Congress remained stalled on resolving the debt ceiling Government policies
20110802 - As fears about a weak US economy were enflamed after investors got another disappointing economic report - this time on consumer spending Economic indicators
20110804 - As fear about the global economy spooked investors Market sentiment
20110808 - As the debt crisis in Europe, lackluster economic news and a downgrade to the US credit rating Europe (Debt crisis), Credit risk
20110809 ++ After the Federal Reserve said it will keep interest rates exceptionally low until 2013 Government policies
20110810 - As investors were confronted with mounting fears about Europe’s ongoing debt crisis, this time in France Europe (Debt crisis)
20110811 ++ On positive earnings and labor market news Economic indicators
20110818 - As renewed concerns about the US and global economies Market sentiment
20110823 ++ Following a report from the FDIC that showed the number of failing banks shrank for the first time in nearly five years Economic indicators
20110829 ++ A Greek bank deal, a solid US consumer spending report and relief that Hurricane Irene caused less damage than expected Government policies
20110902 - After a government report showing no job growth in August stoked fears that the US may be headed into another recession Economic indicators
20110907 ++ The gains came as concerns over Europe’s debt crisis eased and investors geared up for President Obama’s highly anticipated jobs speech Thursday evening Europe (Debt crisis)
20110909 - As bad news out of Europe kept piling up Europe (Debt crisis)
20110921 - Fed disappoints Government policies
20110922 - Fear factor Market sentiment
20110928 - Worries over stalled global growth prospects and concerns that European leaders may not be moving fast enough to solve the region’s debt problems Market sentiment
20110930 - As investors remain worried about the debt crisis in Europe and the outlook for global economic growth Market sentiment
20111003 - With worries about Greece’s solvency still in the spotlight Market sentiment
20111004 ++ Bear market bounce Market sentiment
20111006 ++ On Euro optimism Europe
20111010 ++ As investors cheered a pledge from European leaders to unveil a plan for solving the eurozone’s debt crisis by the end of the month Government policies
20111017 - As worries about Europe’s debt crisis dominated Market sentiment, Europe (Debt crisis)
@@
20111018 ++ Following a report suggesting that Europe’s bailout fund may get a big boost Market sentiment, Government policies
20111025 - Consumer confidence fell sharply to 39.8 in October Market sentiment
@@
20111027 ++ After European Union leaders agreed to expand Europe’s bailout fund and take major losses on Greek bonds Government policies
@@
20111031 - As investors continued to scrutinize the eurozone debt deal Europe (Debt crisis)
@@
20111101 - New fears about the fate of the European rescue plan Market sentiment
20111109 - Italy fears Europe (Debt crisis)
@@
20111123 - As eurozone fears rumbled on and a preliminary report showed that Chinese manufacturing slowed sharply Europe (Debt crisis), China
20111128 ++ On robust Black Friday sales Economic indicators
@@
20111130 ++ After the Federal Reserve said it will work with other central banks to support the global economy Government policies
20111208 - After European Central Bank President Mario Draghi refused to commit to offering broad assistance to troubled eurozone Government policies
@@
20111220 ++ As concerns about the European debt crisis eased and investors welcomed signs of strength in the US housing market Market sentiment, Economic indicators
@@
20120601 - As Europe’s debt crisis remains unresolved and the US economy is showing new signs of distress Europe (Debt crisis), Market sentiment
20120606 ++ As a report on eurozone economic growth revealed no further weakness Economic indicators
20120621 - China’s manufacturing activity continued to drop off in June China
@@
20120629 ++ A deal among European leaders to help struggling eurozone banks buoyed global markets Government policies
@@
20130415 - Following the news of explosions at the Boston Marathon Global events (terrorist attack)
20140203 - After the Institute for Supply Management’s monthly index showed that manufacturing activity last month expanded at its weakest pace since May Economic indicators
20150821 - As global growth concerns accelerated selling pressure to push the Dow and Nasdaq into correction territory Market sentiment
20150824 - Deep fears about China’s economic slowdown China
@@
20150826 ++ The huge rally represents a rebound following six days of dramatic selling that was driven by serious concerns about how China’s slowing economy will impact the rest of the world China
20150827 ++ The U.S. economy grew 3.7% in the second quarter, a very big upward revision than the first official estimate, 2.3% Economic indicators
20150901 - Following more fears about a slowdown in China’s economy China
20150928 - The Consumer Confidence Index dropped, indicating a potential weakening of the US economy Economic indicators
20160107 - As China’s stock market crashed 7% overnight and crude oil plummeted to the lowest level in more than 12 years China, Oil price
20160202 - As renewed declines in oil prices weighed amid mixed reaction to some key earnings reports Oil price
20160212 ++ As European and American bank stocks — as well as oil prices — bounced sharply Company-specific factors, Oil price
20160113 - Low oil prices, as concerns about global economic slowdown Oil price
@@
20160129 ++ A realization that if the US avoids a recession – as most economists think it will – beaten-down stocks could be a good buy Market sentiment
@@
20160624 - British voters chose to leave the European Union Europe (Brexit)
@@
20160627 - British voters chose to leave the European Union Europe (Brexit)
20160909 - As concerns the Federal Reserve might raise interest rates this month loomed following comments made by key Fed officials Market sentiment, Government policies